# MobileLLM-R1: Exploring the Limits of Sub-Billion Language Model Reasoners with Open Training Recipes

### v2

@@ Abstract
The paradigm shift in large language models (LLMs) from instinctive responses to chain-of-thought (CoT) reasoning has fueled two prevailing assumptions: (1) reasoning capabilities only emerge in sufficiently large models, and (2) such capabilities require training on massive datasets. While the first assumption has already been challenged by recent sub-billion-parameter reasoning models such as Qwen3-0.6B and DeepSeek distilled variants, the second remains largely unquestioned. In this work, we revisit the necessity of scaling to extremely large corpora (>10T tokens) for reasoning emergence. By carefully curating and resampling open-source datasets that we identify as beneficial under our designed metrics, we demonstrate that strong reasoning abilities can emerge with far less data. Specifically, we show that only $\sim$2T tokens of high-quality data are sufficient, and pre-training with 4.2T tokens on the dataset resampled from these $\sim$2T tokens, followed by a established post-training procedure, enables the development of MobileLLM-R1, a series of sub-billion-parameter reasoning models that substantially outperform prior models trained on fully open-sourced data. For example, MobileLLM-R1-950M achieves an AIME score of 15.5, compared to just 0.6 for OLMo-2-1.48B and 0.3 for SmolLM-2-1.7B. Remarkably, despite being trained on only 11.7% of the tokens compared to Qwen3’s proprietary 36T-token corpus for pretraining, MobileLLM-R1-950M matches or surpasses Qwen3-0.6B across multiple reasoning benchmarks. To facilitate further research in this direction, we have released the complete training recipe, data sources, data mixing ratio, and model checkpoints, together with the key insights obtained throughout this study.
[…]
@@ 1 Introduction
Large language models (LLMs) such as GPT (Achiam et al., 2023), Qwen (Yang et al., 2025, 2024), and DeepSeek (Guo et al., 2025) have demonstrated remarkable progress in explicit reasoning. Advances have been driven by scaling model size, expanding training data, and applying post-training techniques such as supervised fine-tuning (SFT) and reinforcement learning (RL). Reasoning LLMs are capable of tackling complex problems by following long chains of thought that incorporate reflection, backtracking, and self-validation. At the same time, reasoning traces have evolved from prompt-based chain-of-thought (CoT) in-context learning (Wei et al., 2022) to models explicitly optimized on long reasoning traces to generate multi-step reasoning sequence (Jaech et al., 2024).
@@
Through capability-aware data curation and probing into the latent factors that govern reasoning, we achieve highly token-efficient pretraining compared to prior work. With only 4.2T training tokens, just 11.7% of Qwen’s 36T, our MobileLLM-R1-950M model, matches or surpasses Qwen3-0.6B (Yang et al., 2025) on multiple reasoning benchmarks, placing itself on the Pareto frontier of accuracy–training-token efficiency trade-off curve (Figure 1). Beyond introducing a high-performing small-scale reasoning model, we share both the insights and the pitfalls encountered along the way, offering a first-hand glimpse into the complex yet fascinating mechanisms behind reasoning models.
We make the following contributions toward building the most accurate fully open-source sub-billion–parameter language model reasoners to date from scratch:
@@
2 How do LLMs gain reasoning capability?
Previous work have studied how SFT and RL affect model’s reasoning capability (Wang et al., 2025; Li et al., 2025). However, if people agree that the reasoning capability is "elicited" in the SFT/RL stages, then how does it gain initially when it is trained from scratch? To the best of our knowledge, there is no prior work answering that.
Large Language Models (LLMs) can be fundamentally understood as context-conditioned pattern reconstruction systems, where the central objective is to model the conditional probability distribution of the next token given its preceding context. This perspective frames LLM training as the process of refining a predictive distribution over the vocabulary space: learning, from large-scale data, the statistical regularities that govern natural language.
From this lens, the transition from a general-purpose LLM to a reasoning-specialized model can be interpreted as a systematic shift in token probability mass. That is, reasoning ability emerges not from an entirely new modeling paradigm, but from a redistribution of probability towards reasoning-relevant continuations when presented with certain contexts. Understanding this shift, both qualitatively and quantitatively, provides a principled way to study how reasoning capability is acquired.
We therefore design a complete training curriculum for building a reasoning-capable LLM from scratch, structured to deliberately guide the model through this probability-space transformation. Our approach parallels well-established principles in educational psychology: just as human learners first undergo a broad-based general education before pursuing domain-specific expertise. We found that a solid knowledge foundation is crucial for language models to stabilize their output and ensure good quality in the subsequent reasoning benchmarks.
In the pre-training stage, the model is first exposed to a diverse corpus that grounds it in human language and general world knowledge. During this stage, it also acquires basic mathematical and reasoning capabilities. We will later show that it is crucial to (1) include both types of data and (2) train them jointly: (1) web data provides fundamental linguistic grounding without divergence, while (2) math data develops reasoning capacity. This stage establishes a robust linguistic and conceptual foundation, enabling coherent next-token prediction across a wide range of inputs.
In the subsequent mid-training stage, we strategically shift the data distribution towards reasoning-rich domains such as mathematics, coding, and structured problem solving. This induces a gradual reallocation of probability mass towards reasoning-oriented continuations—analogous to how specialized education shapes cognitive priors in humans.
Finally, supervised fine-tuning (SFT) serves to align the model with human-preferred behaviors, equipping it with instruction-following capabilities and extended sequence handling. This final stage ensures that the reasoning ability acquired during earlier phases is accessible, controllable, and usable in practical interaction settings.
@@ 2 How do LLMs gain reasoning capability?
3 Final Recipe
@@
3.1 Model Architecture
Our model architecture is based on designs from MobileLLM (Liu et al., 2024) and LLaMA3.2 (Dubey et al., 2024). We adopt the LLaMA3.2 tokenizer with a 128k subword vocabulary. We also incorporate QK-norm to mitigate training instabilities in the self-attention block. Following MobileLLM, we adopt weight sharing between the input and output embeddings to improve parameter efficiency. A complete set of architectural specifications is reported in Table 1.
@@ 3.3.3 Post-training
4 Final Results
In the following, we position our trained MobileLLM-R1 within the context of prior state-of-the-art models and compare their performance. We present results for two sets of models: the base models, evaluated after pre-training and mid-training, and the final models after the complete training pipeline. The experimental settings and full training pipeline can be found in Section 3
Base Model Table 6 compares base reasoning models across multiple benchmarks. We group them into fully open-source models (OLMo (OLMo et al., 2024), SmolLM (Allal et al., 2025), MobileLLM-R1), with weights, data, and training recipes available, and partially open-source models (Qwen (Yang et al., 2025), Gemma (Team et al., 2025), LLaMA (Dubey et al., 2024)), which released model weights and partial training procedures. Compared to fully open-source models, MobileLLM-R1 consistently outperforms both OLMo and SmolLM across all parameter scales. For example, at the 140M scale, MobileLLM-R1 achieves 16.3% GSM8K and 15.9% HumanEval, dramatically surpassing SmolLM2-135M (1.8% and 0.0%, respectively). Compared to prior partially open-source models, such as Qwen3-0.6B, MobileLLM-R1 achieves comparable or superior results despite being trained on substantially fewer tokens (4.2T for MobileLLM-R1 vs. 36T for Qwen3). Notably, MobileLLM-R1-950M attains the highest HumanEval score (46.3%) among all sub-1B models, significantly outperforming Qwen3-0.6B (30.5%).
Table 6: Performance comparison of reasoning base models across multiple benchmarks. Here, CommonSense Avg. denotes an average of 8 tasks in CommonSense Reasoning benchmarks including ARC-easy, ARC-challenge, BoolQ, PIQA, SIQA, HellaSwag, OBQA, and WinoGrand. Per-task accuracies are reported in Table 13. Models with fewer than 150M parameters do not yield reliable MMLU scores and are therefore denoted as ‘-’.
Model Size
MATH500
(4-shot, em)
GSM8K
(8-shot, em)
MBPP
(3-shot, pass@1)
HumanEval
(0-shot, pass@1)
CommonSense
(0-shot, avg acc.)
MMLU
(5-shot, acc.)

### v3

# arXiv:2509.24945v3[cs.CL] 27 Feb 2026

@@ Abstract
The paradigm shift in large language models (LLMs) from instinctive responses to chain-of-thought (CoT) reasoning has fueled two prevailing assumptions: (1) reasoning capabilities only emerge in sufficiently large models, and (2) such capabilities require training on massive datasets. While the first assumption has already been challenged by recent sub-billion-parameter reasoning models such as Qwen3-0.6B and DeepSeek distilled variants, the second remains largely unquestioned. In this work, we revisit the necessity of scaling to extremely large corpora (>10T tokens) for reasoning emergence. By carefully curating and resampling open-source datasets that we identify as beneficial under our designed metrics, we demonstrate that strong reasoning abilities can emerge with far less data. Specifically, we show that only $\sim$2T tokens of high-quality data are sufficient, and pre-training with 4.2T tokens on the dataset resampled from these $\sim$2T tokens, followed by a established post-training procedure, enables the development of MobileLLM-R1, a series of sub-billion-parameter reasoning models that substantially outperform prior models trained on fully open-sourced data. For example, MobileLLM-R1-950M achieves an AIME score of 15.5, compared to just 0.6 for OLMo-2-1.48B and 0.3 for SmolLM-2-1.7B. Remarkably, despite being trained on only 11.7% of the tokens compared to Qwen3’s proprietary 36T-token corpus for pretraining, MobileLLM-R1-950M matches or surpasses Qwen3-0.6B across multiple reasoning benchmarks. To facilitate further research in this direction, we have made the models (https://huggingface.co/collections/facebook/mobilellm-r1) and code (https://github.com/facebookresearch/MobileLLM-R1) publicly available, along with the complete training recipe, data sources, and data mixing ratios.
@@ 1 Introduction
Large language models (LLMs) such as GPT (Achiam et al., 2023), Qwen (Yang et al., 2025; 2024), and DeepSeek (Guo et al., 2025) have demonstrated remarkable progress in explicit reasoning. Advances have been driven by scaling model size, expanding training data, and applying post-training techniques such as supervised fine-tuning (SFT) and reinforcement learning (RL). Reasoning LLMs are capable of tackling complex problems by following long chains of thought that incorporate reflection, backtracking, and self-validation. At the same time, reasoning traces have evolved from prompt-based chain-of-thought (CoT) in-context learning (Wei et al., 2022) to models explicitly optimized on long reasoning traces to generate multi-step reasoning sequence (Jaech et al., 2024).
@@
Through capability-aware data curation and probing into the latent factors that govern reasoning, we achieve highly token-efficient pretraining compared to prior work. With only 4.2T training tokens, just 11.7% of Qwen’s 36T, our MobileLLM-R1-950M model, matches or surpasses Qwen3-0.6B Yang et al. (2025) on multiple reasoning benchmarks, placing itself on the Pareto frontier of accuracy–training-token efficiency trade-off curve (Figure 1). Beyond introducing a high-performing small-scale reasoning model, we share both the insights and the pitfalls encountered along the way, offering a first-hand glimpse into the complex yet fascinating mechanisms behind reasoning models.
Our contributions are as follows:
@@
2 Pre-training: Balance of Capabilities
The notion of reasoning in large language models (LLMs) remains both complex and contested. While the term is sometimes used to describe a model’s ability to engage in structured, multi-step inference (Wei et al., 2022; Kojima et al., 2022), it has also become a proxy for improved performance on challenging benchmarks (Srivastava et al., 2022). In this work, we adopt a pragmatic stance: we treat gains on reasoning-centric benchmarks as reasonable evidence of enhanced reasoning behaviors, while remaining cautious about equating such gains with genuine reasoning ability in the cognitive sense (Bender & Koller, 2020). Concretely, this entails selecting informative datasets that most effectively enhance the target capability (Section 2.1) and optimizing their combination ratios to maximize knowledge acquisition within the fixed token budget (Section 2.2). Figure 2 illustrates the training pipeline with the full procedure in Appendix A.
2.1 Selecting Informative Datasets for Target Capability
Hierarchical Rejection Sampling.
To derive a compact capability-probing datasets for each domain, we employ a hierarchical rejection sampling pipeline that integrates multiple classifier- and model-based filters. The objective is to construct a small yet representative target dataset for each capability, such that it can serve as a faithful proxy for reasoning performance while dramatically reducing overall volume during evaluation. For each corpus in Table 5, we first apply the FineWeb-Edu classifier (Penedo et al., 2024) to select samples with high educational value, retaining only those with classifier scores above $4$. Next, we incorporate model-based evaluation by scoring each remaining sample using the Ask-LLM paradigm (Sachdeva et al., 2024). The evaluation prompt asks the model to judge whether a sample should be included in a reasoning-probing dataset, framed as a binary classification task (“1” for inclusion, “0” for exclusion). Rather than relying solely on the hard prediction, we record the probability assigned to “1” as a graded measure of the model’s confidence in the example’s reasoning relevance. For all Ask-LLM scoring, we select the top 10% samples within each dataset. This step complements classifier-based quality filtering by directly capturing signals of reasoning relevance, consistent with recent findings that costly, fine-grained quality samplers can outperform simple maximum-coverage approaches in terms of data efficiency (Sachdeva et al., 2024; Pang et al., 2025; Chen & Zhou, 2025). Next, we apply a domain-specific prompt to Ask-LLM for each capability with specific emphasis on code, math, general knowledge or combined. Finally, we perform semantic deduplication across corpora, shrinking each dataset in Table 5 to a subset of roughly 10,00010{,}000 examples. This yields the representative datasets $\mathcal{D^{R}}_{i}$, each containing highly representative samples for its corresponding corpus.
[…]
2.1.2 Disentangling the Impact of Data Sources
Group Impact via Loss Delta.
[…]
Leave-One-Out Ablations.
We operationalize Eq. 1 by training models under leave-one-out settings and measuring the resulting differences in loss across benchmarks. Together, these analyses highlight not only which sources matter most, but also how much marginal benefit additional data from a given source provides. This methodology allows us to disentangle the contributions of heterogeneous data sources to reasoning-related performance in code, knowledge, and mathematics.
In contrast, domain-specific datasets primarily strengthen their respective domains: Starcoder substantially improves code performance (and, interestingly, math), while math-focused corpora primarily benefit math. However, their transfer to general knowledge is limited. An unexpected observation is that Starcoder benefits math more than OpenWeb-Math benefits code, a reversal of the commonly held view that mathematical data contributes disproportionately to coding ability Lewkowycz et al. (2022). Finally, Wikipedia appears to contribute little to math or code compared to web or domain-specific data, yet remains necessary as a structured and reliable source of factual knowledge.
2.2 Datamixing via Cross-Capability Self-Influence
[…]
Then, the joint influence of a sample is computed as
Each source dataset $g$ is then assigned a sampling weight ($w_{g}$):
3 Mid-training: Knowledge Compression
[…]
(2) Dataset-level influence for adaptive data mixing. Beyond sample-level filtering, we aggregate influence scores to the dataset level, enabling adaptive control of the mixing ratio among mid-training datasets, according to Eqs. 4 and 5).
$\theta_{t+1}=\mathrm{MidTrain}(\theta_{t},\mathcal{D}_{t}).$ (7)
Intuition: Distributional Compression of Influence.
The compression phases can be viewed as iteratively distilling the mid-training dataset in alignment with the model’s evolving capacity throughout training. In early phases, the influence scores are more varied because the model $\theta_{t}$ is still under-trained. However, as $t$ increases, the model becomes better aligned with the target distribution, and its estimates of sample importance are narrowed down (See Figure 5). This recursive interplay produces increasingly refined datasets: uninformative (or negative-influence) samples are discarded, thus amplifying the impact of high-value samples. Conceptually, compression phases mimic an iterative denoising process, where each step sharpens the signal from $\mathcal{D}^{(\mathrm{target})}$ against the noisy background of $\mathcal{D}^{(\mathrm{raw})}$. We terminate the iteration until the distribution of influence converges to approximately zero.
4 Experimental Results
Using the datasets from Sections 2 and 3, we obtain MobileLLM-R1-base. Given that our primary goal is to elucidate how data curation in pre-training and mid-training builds strong small reasoning models, we leverage established supervised fine-tuning (SFT) datasets. We first apply Tülu-3-SFT (Lambert et al., 2024) dataset for instruction alignment and OpenScienceReasoning-2, OpenCodeReasoning-2 (Ahmad et al., 2025) and OpenMathReasoning (Moshkov et al., 2025) for reasoning-oriented SFT to extend context and elicit long chain-of-thought reasoning. We validate our training process and data choices, compare our model against prior work trained on the same reasoning SFT datasets.
Stage 1 Stage 2 MATH GSM8K LCBv6 MMLU
Ablation on Stage 1
w/o Tulu-3 M + C + S 56.2 68.2 13.1 44.0
w/ Tulu-3 M + C + S 57.8 68.5 13.7 43.7
Ablation on Stage 2
Tulu-3 M 57.4 68.2 0.0 43.1
Tulu-3 C 16.2 31.0 12.0 39.9
Tulu-3 S 23.8 62.2 3.4 45.6
Tulu-3 M + C 58.4 65.6 10.9 40.4
Tulu-3 M + S 60.0 66.9 0.6 45.0
Tulu-3 C + S 29.4 65.3 14.3 44.4
Tulu-3 M + C + S 57.8 68.5 13.7 43.7
Joint Ablation
Tulu-3 + (M + C + S) 56.2 53.1 14.9 44.0
Tulu-3 (M + C + S) 57.8 68.5 13.7 44.0
Table 2: Evaluation of reasoning capabilities elicited by different models when fine-tuned on the same reasoning supervised finetuning (SFT) dataset. Baseline models use their instruct checkpoints; our model uses intermediate Tulu3-SFT checkpoints, denoted with *. All models are trained for one epoch on the joint reasoning SFT corpus (OpenMathReasoning + OpenScienceReasoning-2 + OpenCodeReasoning-2).
Model Size MATH GSM8K LCBv6
SmolLM2-135M-Instruct 135M 3.2 1.6 0.6
MobileLLM-R1-140M* 140M 4.8 3.7 1.1
SmolLM2-360M-Instruct 362M 5.2 7.4 3.4
MobileLLM-R1-360M* 359M 19.2 23.8 4.0
OLMo-2-0425-1B-SFT 1.48B 53.0 58.8 11.4
SmolLM2-1.7B-Instruct 1.71B 41.4 50.5 7.4
MobileLLM-R1-950M* 949M 57.8 68.5 13.7
Post-training process ablation: Our ablation studies (Table 2) reveal several key insights into the two-stage post-training pipeline. (1) Instruction-following supervision (Tulu-SFT) provides crucial alignment signals that make subsequent reasoning adaptation significantly more effective than starting directly with reasoning data. (2) Domain-specific reasoning corpora (math, science, code) yield consistent gains on their respective benchmarks, while scientific reasoning data further exhibits strong cross-domain transfer to math and code. (3) Symbolic reasoning improvements often trade off with factual knowledge retention, as introducing math or code data reduces MMLU performance, particularly in smaller models with limited capacity. (4) Decoupling alignment and reasoning proves essential: a staged approach (Tulu first, then reasoning data) consistently outperforms joint training, especially on math and general reasoning benchmarks.
Comparison with baselines on identical reasoning SFT: To disentangle the contribution of curated pre-training and mid-training data from that of high-quality post-training data, we conducted an ablation study. Specifically, we finetune all baseline instruct models, as well as the MobileLLM-R1 general supervised fine-tuned model (trained for 2 epochs on the Tulu dataset), on the joint reasoning SFT corpus (OpenMathReasoning + OpenScienceReasoning-2 + OpenCodeReasoning-2) for one epoch. Our results in Table 2 show that, even under identical supervised fine-tuning, models with stronger pre-training and mid-training exhibit more robustly embedded knowledge, which in turn facilitates the elicitation of reasoning capabilities during post-training. In that sense, MobileLLM-R1 consistently outperforms prior models trained on fully open-source corpora, such as OLMo-2 and SmolLM, on reasoning benchmarks. Notably, our 140M and 360M checkpoints achieve substantial gains over SmolLM baselines, while our 950M model surpasses both OLMo-2 1.48B and SmolLM-1.7B, despite its significantly smaller size.
4.1 Final Results
In the following, we position our trained MobileLLM-R1 within the context of prior state-of-the-art models and compare their performance. We present results for two sets of models: the base models, evaluated after pre-training and mid-training, and the final models after the complete training pipeline. The experimental settings and full training pipeline can be found in Section A
Base Model Figure 8 compares base reasoning models across multiple benchmarks. We group them into fully open-source models (OLMo OLMo et al. (2024), SmolLM Allal et al. (2025), MobileLLM-R1), with weights, data, and training recipes available, and partially open-source models (Qwen Yang et al. (2025), Gemma Team et al. (2025), LLaMA Dubey et al. (2024)), which released model weights and partial training procedures. Compared to fully open-source models, MobileLLM-R1 consistently outperforms both OLMo and SmolLM across all parameter scales. For example, at the 140M scale, MobileLLM-R1 achieves 16.3% GSM8K and 15.9% HumanEval, dramatically surpassing SmolLM2-135M (1.8% and 0.0%, respectively). Compared to prior partially open-source models, such as Qwen3-0.6B, MobileLLM-R1 achieves comparable or superior results despite being trained on substantially fewer tokens (4.2T for MobileLLM-R1 vs. 36T for Qwen3). Notably, MobileLLM-R1-950M attains the highest HumanEval score (46.3%) among all sub-1B models, significantly outperforming Qwen3-0.6B (30.5%).
Post-trained Model Figure 9 presents the performance of post-trained models. Notably, on LiveCodeBench, small models below 400M parameters struggle to produce reliable outputs. In contrast, MobileLLM-R1-360M achieves 5.1 points, surpassing even models with over 1B parameters, such as SmolLM2-1.7B, Gemma3-1B, and LLaMA3.2-1B. Remarkably, MobileLLM-R1-950M demonstrates a substantial accuracy gain over Qwen3-0.6B on LiveCodeBench and even matches the performance of much larger state-of-the-art models, such as DeepSeek-R1-Distill-Qwen-1.5B. Across Math and AIME benchmarks, MobileLLM-R1 consistently outperforms other fully open-source models and achieves scores comparable to the partially open-source Qwen3 series. See Appendix B.1 for detailed comparisons.
5 Related Work
The advent of GPT-3 (Brown et al., 2020) highlighted the transformative potential of large language models (LLMs), spurring research on both proprietary models (e.g., Claude (Anthropic, 2024)) and open-source alternatives (e.g., LLaMA (Touvron et al., 2023a; b; Dubey et al., 2024), Gemma (Team et al., 2024a; b; 2025), Qwen (Team, 2024; Yang et al., 2025)). Scaling laws have been central to understanding how larger models elicit emergent capabilities and potential performance singularities, while efficiency-accuracy trade-offs in smaller models are increasingly studied to optimize computational resources. MobileLLM Liu et al. (2024) pioneered on-device LLM deployment, followed by high-performance small models such as OLMO OLMo et al. (2024) and SmolLM Allal et al. (2025), reflecting a broader trend toward transparency, including open-sourcing weights, training data, and full pipelines.
Research has also shifted from instinctive response to explicit reasoning thinking, exemplified by OpenAI’s O1 (Jaech et al., 2024) and DeepSeek-R1 Guo et al. (2025). Qwen Yang et al. (2025) is another example of a high-quality reasoning model, with smaller variants achieving state-of-the-art benchmark results. Prior studies suggest reasoning emerges only after extremely large-scale pretraining Yang et al. (2025). In contrast, our work shows that a small model can attain strong reasoning abilities using only 4.2T pretraining tokens—comparable to Qwen trained on 36T tokens. We release our full data collection and training pipeline with a detailed rationale for each data selection, ensuring maximal transparency and reproducibility.
6 Conclusion
Ethics Statement
Our work investigates methods for optimizing the training and deployment of small-scale language models. The models and datasets used in this study are publicly available and widely adopted in the research community. We did not collect new human subject data, nor did we rely on sensitive or proprietary sources. Potential risks primarily concern the general risks associated with language models. While these risks are not specific to our contributions, we acknowledge them and emphasize that our focus is methodological rather than on downstream deployment. We encourage future applications of our techniques to carefully assess ethical considerations in their respective domains.
Reproducibility Statement
We follow the ICLR reproducibility guidelines. All datasets used are publicly available, as we revealed in Section A.3. We describe data processing procedures, model architectures, training configurations, and hyperparameters in detail in Sections A. We will release code and trained model checkpoints to support full reproducibility.
Appendix
Appendix A Final Recipe
@@
A.1 Model Architecture
Our model architecture is based on designs from MobileLLM Liu et al. (2024) and LLaMA3.2 Dubey et al. (2024). We adopt the LLaMA3.2 tokenizer with a 128k subword vocabulary. We also incorporate QK-norm to mitigate training instabilities in the self-attention block. Following MobileLLM, we adopt weight sharing between the input and output embeddings to improve parameter efficiency. A complete set of architectural specifications is reported in Table 3.
@@ A.3.3 Post-training
Appendix B Comparison with Previous Methods
B.1 Detailed Accuracy Comparison
Table 8: Performance comparison of reasoning base models across multiple benchmarks. Here, CommonSense Avg. denotes an average of 8 tasks in CommonSense Reasoning benchmarks including ARC-easy, ARC-challenge, BoolQ, PIQA, SIQA, HellaSwag, OBQA, and WinoGrand. Models with fewer than 150M parameters do not yield reliable MMLU scores and are therefore denoted as ‘-’.
Model Size MATH500 (4-shot, em) GSM8K (8-shot, em) MBPP (3-shot, pass@1) HumanEval (0-shot, pass@1) CommonSense (0-shot, avg acc.) MMLU (5-shot, acc.)